# Perturbative soft graviton theorems in de Sitter spacetime

### v2

@@ Abstract
We consider soft graviton scattering for a theory where Einstein’s gravity is minimally coupled to a scalar field in the presence of a cosmological constant, i.e. in a background de Sitter space. Employing a perturbative expansion in a small cosmological constant, we compute the leading, subleading and sub-subleading corrections to Weinberg’s soft graviton amplitude in the static patch of de Sitter space. We observe similar universal features of the soft graviton amplitude as found in Sayali_Diksha for the soft photons.
@@ 1 Introduction
Recently, the soft photon theorems in de Sitter space was studied using the perturbative approach Sayali_Diksha . The perturbative corrections to the flat space soft photon theorem in $d$ dimensions are obtained by analyzing the soft behavior of the perturbative $\mathcal{S}$-matrix defined in the static patch Emil_2021 , and the universality of these corrections was analyzed. This analysis has not yet done for the soft graviton theorems in de Sitter space. It is interesting to see if the universality the soft graviton theorems in de Sitter space is consistent with that of the soft photon theorems. In this paper, we use the same perturbative approach to derive soft graviton theorems in four dimensions in the de Sitter space. We confine scattering processes to a compact region of size $R$ Sayali_Diksha ; Mandal_Banerjee , which is much smaller than the de Sitter curvature length $l$, inside the static patch and define asymptotic states at the boundary of $R$. This approach is different than that of Mao_Zhou , where the asymptotic in- and out-states are defined on the cosmological horizons and soft graviton theorems are obtained via Ward identities.
@@
This paper is organized as follows: In section (2), we briefly describe the geometry of the de Sitter space and define appropriate coordinates for our analysis. We review the scalar fields in the de Sitter background in section (3), which have been studied in details in Sayali_Diksha for generic dimensions $d$. Next, we provide linearized Einstein field equation in the de Sitter background in section (4.1) and obtain mode solutions in section (4.2). By defining the inner product of the modes, we check their orthogonality and consider only the orthogonal set of modes in our analysis. Then, we derive the LSZ formula for gravitons by utilizing the inner product of modes in section (4.3). In section (5), we define the perturbative $\mathcal{S}$-matrix in the static patch of the de Sitter space as in Sayali_Diksha , which differs from the perturbative $\mathcal{S}$-matrix defined in global de Sitter space in Marolf_2013 . Utilizing the LSZ formula, we compute the $\mathcal{S}$-matrix for the scattering process depicted in fig. (2) in section (6). We compute perturbative corrections [^1] By perturbative corrections, we refer to the $\mathcal{O}(l^{-1})$ and $\mathcal{O}(l^{-2})$ corrections to the flat space soft factor, which should not be confused with leading, subleading, and sub-subleading corrections. As analyzed in section (6.2), the $\mathcal{O}(l^{-2})$ corrections include terms of order $1/k^{3}$ and $1/k^{2}$, which do not contribute at the same order. The former is more leading than the latter. Thus, $\mathcal{O}(l^{-2})$ corrections split into subleading and sub-subleading corrections. to the flat space soft factor and deduce leading, subleading, and sub-subleading corrections in the subsequent subsections (6.1) and (6.2). Finally, we conclude by summarizing our results and providing some remarks in section (7).
@@ 2 De Sitter spacetime
in the ambient Minkowski space $\eta_{AB}$, where $X^{A}$ are the Cartesian coordinates and $A$ runs from 00 to $4$. As we shall see, the metric of such a spacetime is conformally equivalent to the flat space. This is best seen as a stereographic projection of the hyperboloid to the flat Minkowski space. To do so, we use the stereographic mapping as given in Aldrovandi_1995 ; Sayali_Diksha .
@@
Where, $\eta_{\mu\nu}$ is the Minkowski metric. The advantage of using the above coordinate system is that the metric is conformally flat. We must note that there is a coordinate singularity at $x^{2}=-4l^{2}$, which however does not affect our analysis. Because, we confine a scattering process to a compact region of size $R$ in the static patch (fig. (1)), where $R<<l$. Therefore, all the points in region $R$ have $x^{\mu}<<l$ for every component “$\mu$”, and hence the above metric takes the following form up to $\mathcal{O}(l^{-2})$:
@@ 3.1 Scalar modes
where $\nabla^{2}$ is the de-Alembertian operator in the de Sitter spacetime and $m$ is the mass of the scalar field $\phi$. Considering corrections up to $\mathcal{O}(l^{-2})$, this equation has the form
@@
where $\Box$ is the flat space de-Alembertian operator. We follow the deduction given in Sayali_Diksha , where the solution for the scalar field equation in embedding space coordinates $X^{A}$ is given by
@@ 3.3 Propagator for the scalar field
The scalar field propagator up to $\mathcal{O}(l^{-2})$ in the de Sitter space has been derived in Sayali_Diksha for arbitrary dimensions $d$. We thus do not run over the full computation, but instead briefly mention the steps required to derive it and then present the final result. The differential equation for the ptopagator can be stated as
@@
It is not hard to check that $D(x,y)$ satisfies a similar equation (16) in the variable $y$. Though the above form of the propagator is not explicitly symmetric in $x$ and $y$, one can brought it to an explicitly symmetric form using integration by parts Sayali_Diksha . In the following section, we discuss the symmetric form of the propagator.
@@ 3.3.1 Symmetric form of Feynman propagator
Here, we state the symmetric form of Feynman propagator, using the mode solutions $g_{p}$. The Feynman propagator in this case can be decomposed into the flat space propagator and corrections of the order $\mathcal{O}(l^{-1})$ and $\mathcal{O}(l^{-2})$, respectively. Thus, we have
@@
Now, we state the expressions of the flat as well as correction terms in $d=4$. This is already explicitly shown in Sayali_Diksha and we just restate it here so that we can use it in our calculations.
${D}^{(0)}$ $=\int\frac{d^{4}p}{(2\pi)^{4}}\frac{e^{ip\cdot(x-y)}}{p^{2}+m^{2}},$
$D^{(1)}$ $=\int\frac{d^{4}p}{(2\pi)^{4}}\frac{e^{ip\cdot(x-y)}}{p^{2}+m^{2}}\ \Big{[}\frac{3p\cdot x}{2lm}+\frac{3p\cdot y}{2lm}+\frac{i(p\cdot x)^{2}}{2lm}-\frac{i(p\cdot y)^{2}}{2lm}+\frac{imx^{2}}{2l}-\frac{imy^{2}}{2l}\Big{]}=0,$
$D^{(2)}$ $=\int\frac{d^{4}p}{(2\pi)^{4}}\frac{e^{ip\cdot(x-y)}}{p^{2}+m^{2}}\ \Big{[}\frac{15(p\cdot x)^{2}}{8l^{2}m^{2}}+\frac{15(p\cdot y)^{2}}{8l^{2}m^{2}}+\frac{3i(p\cdot x)^{2}(p\cdot y)}{4l^{2}m^{2}}+\frac{9(p\cdot x)(p\cdot y)}{4l^{2}m^{2}}$
@@
As is obvious from the expressions, the propagator stated above is manifestly symmetric in $x$ and $y$. A comparison of this propagator with the asymmetric one is already done in Sayali_Diksha and so we do not re-do it here. We would mention however that $\mathcal{O}(l^{-1})$ correction $D^{(1)}$ vanishes. This is because we can express the variable $x$ as $-ie^{-ip\cdot x}\partial_{p}(e^{ip\cdot x})$ and convert it into a $y$ dependent piece by doing an integration by parts in $p$. Then, it is not hard to show that these manipulations render $D^{(1)}$ to be zero. We note this is consistent with the absense of a $1/l$ piece in the propagator derived in the previous section. Another thing we want to point out is that although the $\mathcal{O}(l^{-2})$ correction to the flat space propagator does not seem to match with that in the previous section, we can easily do integration by parts in $p$ and match it thereafter. Thus, all in all, the symmetric form of the Feynman propagator is consistent with its asymmetric form.
@@ 4.1 Linearized Field Equation
To deduce the linearized Einstein field equation in the de Sitter background, the general spacetime metric can be expanded as Date_Hoque_2016 ; Mao_Zhou
@@
$g^{\mu\nu}=\bar{g}^{\mu\nu}-\kappa h^{\mu\nu}+\mathcal{O}(\kappa^{2}).$ (24)
Where, $\kappa=\sqrt{32\pi G}$ and $\bar{g}_{\mu\nu}$ is the background de Sitter metric. The indices of $h_{\mu\nu}$ are raised and lowered by $\bar{g}_{\mu\nu}$.
@@
where $\bar{\Gamma}^{\alpha}_{\mu\nu}$ and ¯$\bar{\nabla}_{\alpha}$ are respectively the Christoffel symbols and covariant derivatives compatible with the de Sitter metric. Now, we determine the Riemann curvature tensor up to $\mathcal{O}(\kappa)$:
$R_{\mu\nu\alpha}^{\ \ \ \ \beta}=\bar{R}_{\mu\nu\alpha}^{\ \ \ \ \beta}\ +\dfrac{\kappa}{2}\bar{\nabla}_{\nu}\left(\bar{\nabla}_{\mu}h^{\beta}_{\alpha}+\bar{\nabla}_{\alpha}h^{\beta}_{\mu}-\bar{\nabla}^{\beta}h_{\mu\alpha}\right)-\dfrac{\kappa}{2}\bar{\nabla}_{\mu}\left(\bar{\nabla}_{\nu}h^{\beta}_{\alpha}+\bar{\nabla}_{\alpha}h^{\beta}_{\nu}-\bar{\nabla}^{\beta}h_{\nu\alpha}\right),$ (26)
where $\bar{R}_{\mu\nu\alpha}^{\ \ \ \ \beta}$ is the Riemann curvature tensor for the de Sitter space. To obtain the Einstein tensor, we require the Ricci tensor and the Ricci scalar. Contracting the indices $\nu$ and $\beta$ in (26) yields the Ricci tensor
$R_{\mu\alpha}=\bar{R}_{\mu\alpha}+\dfrac{\kappa}{2}\bar{\nabla}_{\nu}\left(\bar{\nabla}_{\mu}h^{\nu}_{\alpha}+\bar{\nabla}_{\alpha}h^{\nu}_{\mu}-\bar{\nabla}^{\nu}h_{\mu\alpha}\right)-\dfrac{\kappa}{2}\bar{\nabla}_{\mu}\bar{\nabla}_{\alpha}h,$ (27)
@@
$R=\bar{R}-\kappa h^{\mu\alpha}\bar{R}_{\mu\alpha}+\dfrac{\kappa}{2}\bar{\nabla}_{\nu}\left(\bar{\nabla}_{\mu}h^{\mu\nu}+\bar{\nabla}^{\mu}h^{\nu}_{\mu}-\bar{\nabla}^{\nu}h\right)-\dfrac{\kappa}{2}\bar{\nabla}_{\mu}\bar{\nabla}^{\mu}h.$ (28)
Let us now consider the vacuum Einstein field equation with a cosmological term.
$E_{\mu\nu}=R_{\mu\nu}-\dfrac{1}{2}g_{\mu\nu}R+\Lambda g_{\mu\nu}=0,$ (29)
@@
$E_{\mu\nu}=\dfrac{1}{2}\left(\bar{\nabla}_{\tau}\bar{\nabla}_{\mu}h^{\tau}_{\nu}+\bar{\nabla}_{\tau}\bar{\nabla}_{\nu}h^{\tau}_{\mu}-\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}h-\bar{\nabla}^{2}h_{\mu\nu}\right)\\ -\dfrac{1}{2}\bar{g}_{\mu\nu}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}h^{\alpha\beta}-\bar{\nabla}^{2}h\right)-\dfrac{3}{l^{2}}h_{\mu\nu}+\dfrac{3}{2l^{2}}\bar{g}_{\mu\nu}h,$ (30)
where the following relations for the de Sitter space are used: $\bar{R}_{\mu\nu}=\dfrac{3}{l^{2}}\bar{g}_{\mu\nu}$, $\bar{R}=\dfrac{12}{l^{2}}$.
@@
Applying the de Donder like gauge condition, ¯$\bar{\nabla}_{\mu}h^{\mu\nu}=\dfrac{1}{2}\bar{\nabla}^{\nu}h$, the linearized field equation reduces to
$E_{\mu\nu}=\left(\bar{\nabla}^{2}-\dfrac{2}{l^{2}}\right)h_{\mu\nu}=0.$ (32)

### v3

@@ Abstract
We consider soft graviton scattering for a theory where Einstein’s gravity is minimally coupled to a scalar field in the presence of a cosmological constant, i.e. in a background de Sitter space. Employing a perturbative expansion in a small cosmological constant, we compute leading, subleading and sub-subleading corrections to Weinberg’s soft graviton amplitude for the tree-level scatterings in the static patch of de Sitter space. We observe similar universal features of the soft graviton amplitude as found in Sayali_Diksha for the soft photons.
@@ 1 Introduction
Recently, the soft photon theorems in de Sitter space was studied using the perturbative approach Sayali_Diksha . The perturbative corrections to the flat space soft photon theorem in $d$ dimensions are obtained by analyzing the soft behavior of the perturbative $\mathcal{S}$-matrix defined in the static patch Emil_2021 , and the universality of these corrections was analyzed. This analysis has not yet been done for the soft graviton theorems in de Sitter space. It is interesting to see if the universality of the soft graviton theorems in de Sitter space is consistent with that of the soft photon theorems. In this paper, we use the same perturbative approach to derive soft graviton theorems in four dimensions in the de Sitter space. Our setup is as follows: We confine scattering processes to a compact region of size $R$ inside the static patch, which is much smaller than the de Sitter curvature length $l$ Sayali_Diksha ; Mandal_Banerjee so that gravitational effects can be treated perturbatively, and much larger than the Planck length so that the gravitational interactions become weak as particles approach the boundaries. The graviton detector is also placed in the same region. The boundaries of this region can be reached in a finite time which is assumed to be much larger than the interaction time scale. Thus, the particles are asymptotically decoupled Hijano_Neuenfeld , which means they are far away from each other and non-interacting at early and late times. Thus, the early and late time Cauchy surfaces are the Hilbert spaces of the incoming and outgoing particles, respectively. Our results are valid in the regime $\sqrt{G}<<R<<l$. Our approach is different than that of Mao_Zhou , where the asymptotic in- and out-states are defined on the cosmological horizons and soft graviton theorems are obtained via Ward identities.
@@
This paper is organized as follows: In section (2), we briefly describe the geometry of the de Sitter space and define appropriate coordinates for our analysis. We review the scalar fields in the de Sitter background in section (3), which have been studied in details in Sayali_Diksha for generic dimensions $d$. Next, we provide linearized Einstein field equation in the de Sitter background in section (4.1) and obtain mode solutions in section (4.2). By defining the inner product of the modes, we check their orthogonality and consider only the orthogonal set of modes in our analysis. Then, we derive the LSZ formula for gravitons by utilizing the inner product of modes in section (4.3). In section (5), we define the perturbative $\mathcal{S}$-matrix in the static patch of the de Sitter space as in Sayali_Diksha , which differs from the perturbative $\mathcal{S}$-matrix defined in global de Sitter space in Marolf_2013 . Utilizing the LSZ formula, we compute the $\mathcal{S}$-matrix for the scattering process depicted in Fig. (2) in section (6). We compute perturbative corrections [^1] By perturbative corrections, we refer to the $\mathcal{O}(l^{-1})$ and $\mathcal{O}(l^{-2})$ corrections to the flat space soft factor, which should not be confused with leading, subleading, and sub-subleading corrections. As analyzed in section (6.2), the $\mathcal{O}(l^{-2})$ corrections include terms of order $1/k^{3}$ and $1/k^{2}$, which do not contribute at the same order. The former is more leading than the latter. Thus, $\mathcal{O}(l^{-2})$ corrections split into subleading and sub-subleading corrections. to the flat space soft factor and deduce leading, subleading, and sub-subleading corrections in the subsequent subsections (6.1) and (6.2). In (6.3), we compute contributions from the second diagram (Fig. (3)) where a soft graviton is attached to an internal line. We do consistency checks of our results in (6.4). Finally, we conclude by summarizing our results and providing some remarks in section (7).
@@ 2 De Sitter spacetime
in the ambient Minkowski space $\eta_{AB}$, where $X^{A}$ are the Cartesian coordinates and $A$ runs from 0 to $4$. As we shall see, the metric of such a spacetime is conformally equivalent to the flat space. This is best seen as a stereographic projection of the hyperboloid to the flat Minkowski space. To do so, we use the stereographic mapping as given in Aldrovandi_1995 ; Sayali_Diksha .
@@
Where, $\eta_{\mu\nu}$ is the Minkowski metric. The advantage of using the above coordinate system is that the metric is conformally flat. We must note that there is a coordinate singularity at $x^{2}=-4l^{2}$, which however does not affect our analysis. Because, we confine a scattering process to a compact region of size $R$ in the static patch (Fig. (1)), where $R<<l$. Therefore, all the points in region $R$ have $x^{\mu}<<l$ for every component $\mu$”, and hence the above metric takes the following form up to $\mathcal{O}(l^{-2})$:
@@ 3.1 Scalar modes
where $\nabla^{2}$ is the d’Alembertian operator in the de Sitter spacetime and $m$ is the mass of the scalar field $\phi$. Considering corrections up to $\mathcal{O}(l^{-2})$, this equation has the form
@@
where $\Box$ is the flat space d’Alembertian operator. We follow the deduction given in Sayali_Diksha , where the solution for the scalar field equation in embedding space coordinates $X^{A}$ is given by
@@ 3.3 Propagator for the scalar field
The scalar field propagator up to $\mathcal{O}(l^{-2})$ in the de Sitter space has been derived in Sayali_Diksha for arbitrary dimensions $d$. We thus do not run over the full computation, but instead briefly mention the steps required to derive it and then present the final result. The differential equation for the propagator can be stated as
@@
It is not hard to check that $D(x,y)$ also satisfies Eq. (16) when $x$ and $y$ are interchanged. Though the above form of the propagator is not explicitly symmetric in $x$ and $y$, one can bring it to an explicitly symmetric form using integration by parts Sayali_Diksha . In the following section, we discuss the symmetric form of the propagator.
@@ 3.3.1 Symmetric form of Feynman propagator
We can write the symmetric form of the Feynman propagator using the orthogonal set of modes $g_{p}$ as follows: Sayali_Diksha
$D(x,y)=-i\int\dfrac{d^{4}p}{(2\pi)^{4}}\dfrac{g_{p}(x)g^{*}_{p}(y)}{p^{2}+m^{2}-i\epsilon}.$ (21)
Substituting Eq. (10) into the above expression, and keeping terms up to $\mathcal{O}(l^{-2})$, the Feynman propagator can be decomposed into the flat space propagator and corrections of the order $\mathcal{O}(l^{-1})$ and $\mathcal{O}(l^{-2})$, respectively. Thus, we have
@@
Now, we state the expressions of the flat as well as correction terms in $d=4$. This has already been explicitly shown in Sayali_Diksha and we just restate it here so that we can use it in our calculations.
${D}^{(0)}$ $=-i\int\frac{d^{4}p}{(2\pi)^{4}}\frac{e^{ip\cdot(x-y)}}{p^{2}+m^{2}},$
$D^{(1)}$ $=-i\int\frac{d^{4}p}{(2\pi)^{4}}\frac{e^{ip\cdot(x-y)}}{p^{2}+m^{2}}\ \Big[\frac{3p\cdot x}{2lm}+\frac{3p\cdot y}{2lm}+\frac{i(p\cdot x)^{2}}{2lm}-\frac{i(p\cdot y)^{2}}{2lm}+\frac{imx^{2}}{2l}-\frac{imy^{2}}{2l}\Big]=0,$
$D^{(2)}$ $=-i\int\frac{d^{4}p}{(2\pi)^{4}}\frac{e^{ip\cdot(x-y)}}{p^{2}+m^{2}}\ \Big[\frac{15(p\cdot x)^{2}}{8l^{2}m^{2}}+\frac{15(p\cdot y)^{2}}{8l^{2}m^{2}}+\frac{3i(p\cdot x)^{2}(p\cdot y)}{4l^{2}m^{2}}+\frac{9(p\cdot x)(p\cdot y)}{4l^{2}m^{2}}$
@@
As is obvious from the expressions, the propagator stated above is manifestly symmetric in $x$ and $y$. A comparison of this propagator with the asymmetric one is already done in Sayali_Diksha and so we do not re-do it here. We would mention however that $\mathcal{O}(l^{-1})$ correction $D^{(1)}$ vanishes. This is because we can express the variable $x$ as $i\partial_{p}$ and convert it into a $y$ dependent piece by doing an integration by parts in $p$. Then, it is not hard to show that these manipulations render $D^{(1)}$ to be zero. We note this is consistent with the absence of a $1/l$ piece in the asymmetric form of the propagator derived in the previous section. Another thing we want to point out is that although the $\mathcal{O}(l^{-2})$ correction to the flat space propagator does not seem to match with that in the previous section, we can easily do integration by parts in $p$ and match it thereafter. Thus, all in all, the symmetric form of the Feynman propagator is consistent with its asymmetric form. [^2] It is worthwhile to mention that the consistency of the two forms of the propagator means one form can be transformed into another by doing integration by parts in $p$.
@@ 4.1 Linearized field equation
To deduce the linearized Einstein field equation in the de Sitter background, a general spacetime metric can be expanded as Date_Hoque_2016 ; Mao_Zhou
@@
$g^{\mu\nu}=\bar{g}^{\mu\nu}-\kappa h^{\mu\nu}+\mathcal{O}(\kappa^{2}),$ (25)
where $\kappa=\sqrt{32\pi G}$, and $\bar{g}_{\mu\nu}$ is the background de Sitter metric. The indices of $h_{\mu\nu}$ are raised and lowered by $\bar{g}_{\mu\nu}$.
@@
where $\bar{\Gamma}^{\alpha}_{\mu\nu}$ and $\bar{\nabla}_{\alpha}$ are respectively the Christoffel symbols and covariant derivatives compatible with the de Sitter metric. Now, we determine the Riemann curvature tensor up to $\mathcal{O}(\kappa)$:
$\mathcal{R}_{\mu\nu\alpha}^{\ \ \ \ \beta}=\bar{\mathcal{R}}_{\mu\nu\alpha}^{\ \ \ \ \beta}\ +\dfrac{\kappa}{2}\bar{\nabla}_{\nu}\left(\bar{\nabla}_{\mu}h^{\beta}_{\alpha}+\bar{\nabla}_{\alpha}h^{\beta}_{\mu}-\bar{\nabla}^{\beta}h_{\mu\alpha}\right)-\dfrac{\kappa}{2}\bar{\nabla}_{\mu}\left(\bar{\nabla}_{\nu}h^{\beta}_{\alpha}+\bar{\nabla}_{\alpha}h^{\beta}_{\nu}-\bar{\nabla}^{\beta}h_{\nu\alpha}\right),$ (27)
where $\bar{\mathcal{R}}_{\mu\nu\alpha}^{\ \ \ \ \beta}$ is the Riemann curvature tensor of the de Sitter space. To obtain the Einstein tensor, we require the Ricci tensor and the Ricci scalar. Contracting the indices $\nu$ and $\beta$ in (27) yields the Ricci tensor
$\mathcal{R}_{\mu\alpha}=\bar{\mathcal{R}}_{\mu\alpha}+\dfrac{\kappa}{2}\bar{\nabla}_{\nu}\left(\bar{\nabla}_{\mu}h^{\nu}_{\alpha}+\bar{\nabla}_{\alpha}h^{\nu}_{\mu}-\bar{\nabla}^{\nu}h_{\mu\alpha}\right)-\dfrac{\kappa}{2}\bar{\nabla}_{\mu}\bar{\nabla}_{\alpha}h,$ (28)
@@
$\mathcal{R}=\bar{\mathcal{R}}-\kappa h^{\mu\alpha}\bar{\mathcal{R}}_{\mu\alpha}+\dfrac{\kappa}{2}\bar{\nabla}_{\nu}\left(\bar{\nabla}_{\mu}h^{\mu\nu}+\bar{\nabla}^{\mu}h^{\nu}_{\mu}-\bar{\nabla}^{\nu}h\right)-\dfrac{\kappa}{2}\bar{\nabla}_{\mu}\bar{\nabla}^{\mu}h.$ (29)
Let us now consider the vacuum Einstein field equation with the cosmological constant term.
$E_{\mu\nu}=\mathcal{R}_{\mu\nu}-\dfrac{1}{2}g_{\mu\nu}\mathcal{R}+\Lambda g_{\mu\nu}=0,$ (30)
@@
$E_{\mu\nu}=\dfrac{1}{2}\left(\bar{\nabla}_{\tau}\bar{\nabla}_{\mu}h^{\tau}_{\nu}+\bar{\nabla}_{\tau}\bar{\nabla}_{\nu}h^{\tau}_{\mu}-\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}h-\bar{\nabla}^{2}h_{\mu\nu}\right)\\ -\dfrac{1}{2}\bar{g}_{\mu\nu}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}h^{\alpha\beta}-\bar{\nabla}^{2}h\right)-\dfrac{3}{l^{2}}h_{\mu\nu}+\dfrac{3}{2l^{2}}\bar{g}_{\mu\nu}h=0,$ (31)
where the following relations for the de Sitter space are used: $\bar{\mathcal{R}}_{\mu\nu}=\dfrac{3}{l^{2}}\bar{g}_{\mu\nu}$, $\bar{\mathcal{R}}=\dfrac{12}{l^{2}}$.
@@
Applying the de Donder gauge condition, $\bar{\nabla}_{\mu}h^{\mu\nu}=\dfrac{1}{2}\bar{\nabla}^{\nu}h$, the linearized field equation reduces to
$\bar{\nabla}^{2}h_{\mu\nu}-\dfrac{1}{2}\bar{g}_{\mu\nu}\bar{\nabla}^{2}h-\dfrac{2}{l^{2}}h_{\mu\nu}-\dfrac{1}{l^{2}}\bar{g}_{\mu\nu}h=0.$ (33)
We can fix the gauge $h=0$ as in the flat space as follows Higuchi . First, we take the trace of the above expression with $h_{\mu\nu}$ replaced by $h^{\prime}_{\mu\nu}$ as
$\left(\bar{\nabla}^{2}+\dfrac{6}{l^{2}}\right)h^{\prime}=0.$ (34)
Let us now define
$h_{\mu\nu}=h^{\prime}_{\mu\nu}+\dfrac{1}{6/l^{2}}\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}h^{\prime},$ (35)
which is traceless from Eq. (34), and satisfies the de Donder gauge condition. Thus, the linearized field equation reduces to
$\left(\bar{\nabla}^{2}-\dfrac{2}{l^{2}}\right)h_{\mu\nu}=0,$ (36)
and the de Donder gauge reduces to the transverse-traceless gauge
$\bar{\nabla}_{\mu}h^{\mu\nu}=0=h.$ (37)
In the following sections, we will drop the bar notations for the de Sitter background.